OpenStudy (anonymous):

what is the value of "y" in the system of the following equations : x=y+3 2x+3y=-4

hartnn (hartnn):

substitute x=y+3 in 2nd equation, what u get ?

OpenStudy (anonymous):

3x-3y=9 (first eq multiplied by 3) 2x+3y=-4 adding the above two eq we have 5x=5 or x=1 now y=x-3 hence y=1-3=-2

OpenStudy (anonymous):

If you substitute the x=y+3 into 2x+3y=-4, you get: 2(y+3) + 3y= -4 Multiply it out to get: 2y+6+3y=-4. Then, combine like terms. 2y+3y= 5y +6=-4 Now you have to solve this like a regular equation: First, subtract 6 from bothsides to get: 5y+6=-4 -6 -6 -------------- 5y= -10 Divide 5 --------------- y= -2 Questions?

OpenStudy (esraa):

y=-2